# From Context to EDUs: Faithful and Structured Context Compression via Elementary Discourse Unit Decomposition

### v2

@@ 3.1 Evaluation of Structural Integrity and Compression
The absence of public benchmarks for fine-grained document structure analysis motivated us to construct a specific dataset named StructBench. We compiled a test set of 248 documents, covering diverse formats (Web pages, PDFs), languages (Chinese, English), and genres (e.g., government files, institutional reports, academic papers, and technical tutorials). Document lengths vary significantly, ranging from 300 to 50,000 words. To ensure high-quality ground truth, documents were parsed, sentence-segmented, and manually annotated for discourse structure by human experts. To enable fair comparison with baselines that may struggle with leaf-level details, we extracted the structural backbone (top-level hierarchy) from the annotations to serve as the ground truth. [^2] The dataset is available at https://huggingface.co/datasets/deeplang-ai/StructBench
@@
We employ two complementary metrics to evaluate structural fidelity:
  • Tree Edit Distance (TED) (ted_zhang1989simple): A micro-level metric that measures structural dissimilarity by computing the minimum number of edit operations (insertion, deletion, substitution) required to transform the predicted tree into the ground truth. Lower TED indicates more precise structural alignment.
  • Document Level Accuracy (DLA): A macro-level metric defined as $\mathrm{DLA}=\frac{\lvert D_{\text{match}}\rvert}{\lvert D_{\text{all}}\rvert}$, where $D_{\text{match}}$ represents the count of documents where the decomposed structure (backbone) perfectly matches the ground truth. This is a rigorous metric requiring zero structural errors.
@@
We compare our LingoEDU (built on Qwen3-4B) against two categories of strong baselines. For all LLMs, we designed specific prompts to instruct them to output hierarchical JSON/Markdown structures:
  • Frontier LLMs: We evaluated SOTA models including GPT-4o/4.1 (gpt-4o), OpenAI o3/o4-mini (o3), Claude 3.7 Sonnet/4 (Claude4), DeepSeek-V3/R1 (deepseekai2025deepseekv3technicalreport), and Qwen3-235B (qwen3). All LLM results are averaged over three distinct runs.
  • Commercial Parsing APIs: We selected Jina Reader [^3] https://jina.ai/reader and Firecrawl [^4] https://www.firecrawl.dev, which are widely used for web-to-markdown conversion. We deployed test documents on a static server to allow URL-based access.
@@
Our method utilizes Qwen3-4B as the backbone. The training involved a two-stage process: (1) Continued Pre-training on $\sim$100k synthetic samples to learn layout patterns, followed by (2) Supervised Fine-Tuning (SFT) on thousands of meticulously manually annotated documents to align with human intent. All experiments were conducted on a Linux operating system running on a high-performance server equipped with an Intel Xeon 2.3GHz CPU, 1960GB of memory, and 8 NVIDIA A800 GPUs, each with 80 GB of VRAM.
Table 1: Performance comparison on StructBench. $\ast$ indicates the model is accessed via API. $\dagger$ denotes local deployment for latency testing. Best results are bolded, and second-best are underlined. Our method achieves SOTA structural accuracy with significantly lower costs.
[…]
GPT-4o General LLM 8.53 36.29% 0.0210 -
GPT-4.1 9.14 37.90% 0.0168 -
OpenAI o3 8.01 35.48% 0.0168 -
OpenAI o4-mini 8.45 36.29% 0.0092 -
Claude-3.7-Sonnet 9.98 35.48% 0.0286 -
Claude-4 7.98 41.53% 0.0286 -
Gemini-2.5-flash 8.12 33.74% 0.0040 -
Gemini-2.5-pro 8.15 35.89% 0.0162 -
DeepSeek-V3 9.12 34.68% 0.0012 -
DeepSeek-R1 8.44 35.08% 0.0046 -
Qwen3-32B 8.55 34.01% 0.0012 10.17
Qwen3-235B 9.81 27.02% 0.0012 -
Jina-Reader Parser API 17.04 - 0.0004 -
Firecrawl 16.81 - 0.0007 -
Our Method (LingoEDU) Specialized 5.67 46.77% 0.0007 1.20
@@
Table 1 highlights the superiority of our explicit training paradigm.
  • General LLMs struggle with precision: Even advanced reasoning models like OpenAI o3 and DeepSeek-R1 plateau around a TED of 8.0-9.0. Qualitative analysis reveals they often hallucinate non-existent sub-sections or flatten deep hierarchies to save generation tokens.
  • Parsing APIs lack semantic depth: While low-cost, Jina and Firecrawl exhibit high TED scores (>16), indicating they rely on shallow HTML tags and fail to capture implicit discourse structures found in complex PDFs.
  • Specialized Efficiency: Our LingoEDU Decomposer achieves a remarkable TED of 5.67 and DLA of 46.77%, significantly outperforming Claude-4 (+5.24% absolute DLA). This confirms that structural understanding requires dedicated supervision rather than emergent prompting.
@@
Ablation Studies.
Table 2 validates the effectiveness of our design. First, the performance gap between “Indices Only” and our method highlights that explicit text generation acts as a crucial semantic anchor. Second, the model exhibits strong data efficiency, maintaining over 94% of its peak performance even when scaled down to 20% of the training data.
@@
Ablation Setting Variant DLA (%)
Output Formulation Indices Only 33.47
Indices + Text (Ours) 46.77
Data Scaling 20% Data 44.35
50% Data 45.97
100% Data 46.77
@@ 3.2 Main Results on Downstream Long-Context Tasks
In this section, we address the second research question: Does structure-aware compression tangibly enhance performance for downstream tasks? We evaluate the utility of our EDU-based Context Compressor across two distinct scenarios: standard long-context benchmarks and complex Deep Search pipelines.
@@ 3.2.1 General Long-Context Understanding
We first evaluate our method on LongBench (bai-etal-2024-longbench), a comprehensive benchmark covering multi-document QA, summarization, and few-shot learning.
@@
We compare three settings:
  1. 1.
    Standard: Feeds the full (or truncated) original context directly to the LLM.
  2. 2.
    Self-Sum: Uses the LLM itself to generate abstractive summaries of the context before processing.
  3. 3.
    Glyph(cheng2025glyph): Applies an implicit compression baseline (where applicable) serving as a reference for latent-based methods.
@@
Table 3 compares performance across Gemini-2.5-Pro and GPT-4.1 backbones. Overall, our EDU-based approach yields consistent improvements over the Standard baseline, with relative gains ($\Delta$) peaking at +23.66% on few-shot tasks.
@@
  • Multi-Document QA (HotpotQA, Musique): Our method outperforms Self-Sum (e.g., +2.22 points on HotpotQA with GPT-4.1). Abstractive methods often lose critical entities required for multi-hop reasoning. In contrast, our explicit tree structure preserves precise evidence chains via original text indices, enabling the model to “look up” exact details without hallucination.
  • Summarization Tasks: While Self-Sum is naturally strong here, our method remains competitive (e.g., surpassing Self-Sum on MultiNews by +1.44 points). Notably, it significantly outperforms the external Glyph baseline, suggesting that retaining hierarchical structure is more effective than latent compression for information retention.
Table 3: Performance comparison on LongBench. Standard denotes the baseline without compression, Self-Sum utilizes LLM’s built-in summarization, and Ours (LingoEDU) represents our structure-aware compression. $\Delta$ indicates the relative improvement of Ours over the Standard baseline. Bold values indicate best performance; underlined values indicate second-best.
Task Type Dataset Glyph Gemini-2.5-Pro GPT-4.1
Standard Self-Sum Ours (LingoEDU) $\Delta$ Standard Self-Sum Ours (LingoEDU) $\Delta$
Multi-Doc QA HotpotQA 66.42 35.20 37.78 40.46 +14.94% 65.83 67.89 70.11 +6.50%
2WikiMQA 72.98 38.10 39.90 40.91 +7.38% 72.98 74.39 74.68 +2.33%
Musique - 28.55 30.77 31.22 +9.35% 51.90 53.48 54.86 +5.70%
DuReader - 7.15 7.79 8.12 +7.69% 21.80 23.51 25.34 +16.24%
Summarization GovReport 25.53 4.10 4.34 4.25 +2.44% 29.97 30.98 31.56 +2.94%
QMSum 19.78 15.80 16.53 16.17 +2.34% 22.84 22.53 23.30 +0.61%
MultiNews - 4.05 4.44 4.85 +19.75% 20.85 22.06 23.50 +5.80%
VCSum - 5.80 6.17 6.36 +9.66% 12.50 13.71 14.62 +8.96%
Few-shot TREC 82.62 46.50 49.00 57.50 +23.66% 77.00 80.50 80.00 +3.90%
TriviaQA 88.54 59.85 62.31 63.25 +1.25% 90.07 93.69 93.76 +4.10%
SAMSum - 20.45 21.89 23.80 +11.39% 39.20 40.79 41.68 +6.33%
LSHT - 26.10 29.50 35.48 +3.45% 48.60 50.50 52.50 +8.02%
@@ 3.2.2 Impact on Deep Search
Unlike standard retrieval, Deep Search involves aggregating information from diverse, often noisy web sources to answer complex, open-ended queries. We integrated the LingoEDU module into a Deep Search pipeline and evaluated it on two challenging benchmarks designed to push the limits of current LLMs:
@@
Deep search engines frequently ingest raw web pages cluttered with ads, navigation bars, and irrelevant links. As shown in Table 4, on the noise-intensive BrowseComp-ZH, our method delivers dramatic gains. Specifically, Qwen3-235B and DeepSeek-V3.1 show relative improvements of 50.59% and 32.53%, respectively. This confirms that our Decomposer acts as a semantic filter: by identifying logical EDUs, it effectively prunes “structural dead branches” while preserving the core content, a capability that standard summarization lacks in such high-entropy Chinese web contexts.
@@
Critically, our method remains additive even for the most advanced models like GPT-5 and Gemini 3 Pro (Table 4, rows 4-6). This indicates that even as model capacity grows, handling unstructured noise and long-context reasoning remain bottlenecks. Our explicit compression provides a complementary signal that enhances the internal processing of SOTA LLMs.
Table 4: Ablation study of the LingoEDU module on Deep Search benchmarks. Accuracy scores (%) are reported. Base: Standard Deep Search without compression; Self-Sum: Query-focused summarization; Ours (LingoEDU): Structural decomposition. $\Delta$ denotes the relative improvement of Ours over the Base baseline. Bold indicates best performance per row.
Model Backbone HLE (Academic Reasoning) BrowseComp-ZH (Noisy Web)
DeepSeek-R1 9.0 9.5 13.6 +51.11% 18.8 19.4 20.4 +8.51%
Qwen3-235B-Thinking 14.2 14.7 15.5 +9.15% 8.5 9.0 12.8 +50.59%
DeepSeek-V3.1 14.5 14.8 15.6 +7.59% 29.2 29.7 38.7 +32.53%
Closed-Source Models
GPT-5 25.0 25.9 27.1 +8.40% 29.0 29.8 31.8 +9.66%
Claude Opus 4.1 14.0 14.8 15.5 +10.71% 20.8 21.5 23.2 +11.54%
Gemini 3 Pro 26.1 26.7 30.1 +15.33% 47.5 48.0 49.0 +3.16%

### v3

@@ 3.1.1 Experiment Settings
The absence of public benchmarks for fine-grained document structure analysis motivated us to construct a specific dataset named StructBench. We compiled a test set of 248 documents, covering diverse formats (Web pages, PDFs), languages (Chinese, English), and genres (e.g., government files, institutional reports, academic papers, and technical tutorials). The dataset spans 10 distinct genres, primarily focusing on complex structures such as academic papers, government files, business reports, and technical tutorials (full distribution in Appendix B). Document lengths vary significantly, ranging from 300 to 50,000 words. To ensure high-quality ground truth, documents were parsed, sentence-segmented, and manually annotated for discourse structure by human experts. To enable fair comparison with baselines that may struggle with leaf-level details, we extracted the structural backbone (top-level hierarchy) from the annotations to serve as the labels.  [^2] The dataset is available at https://huggingface.co/datasets/deeplang-ai/StructBench
[…]
GPT-4o General LLM 6.22 29.03% 5.21 -
GPT-4.1 6.35 37.90% 4.17 -
OpenAI o3 5.51 28.63% 4.17 -
OpenAI o4-mini 5.87 32.66% 2.28 -
Claude-3.7-Sonnet 6.65 35.08% 7.09 -
Claude-4-Sonnet 5.08 43.15% 7.09 -
Gemini-2.5-Flash 5.82 27.82% 0.99 -
Gemini-2.5-Pro 5.61 32.66% 4.02 -
DeepSeek-V3 6.32 33.47% 0.30 -
DeepSeek-R1 6.26 30.65% 1.14 -
Qwen3-32B 9.49 24.90% 0.26 10.17
Qwen3-235B 9.93 17.89% 0.11 -
Jina-Reader Parser API 17.04 - 0.10 -
Firecrawl 16.81 - 0.17 -
Our Method (LingoEDU) Specialized 4.77 49.60% 0.17 1.20
Table 1: Performance comparison on StructBench. $\ast$ indicates the model is accessed via API. $\dagger$ denotes local deployment for latency testing using equivalent hardware. Costs are calculated for the entire test set (approx. 248 documents). Best results are bolded, and second-best are underlined.
@@
We employ two complementary metrics to evaluate structural fidelity. Tree Edit Distance (TED) (ted_zhang1989simple) acts as a micro-level metric to measure structural dissimilarity by computing the minimum number of edit operations (insertion, deletion, substitution) required to transform the predicted tree into the ground truth, where a lower TED indicates more precise structural alignment. Complementarily, Document Level Accuracy (DLA) serves as a macro-level metric defined as $\mathrm{DLA}=\frac{\lvert D_{\text{match}}\rvert}{\lvert D_{\text{all}}\rvert}$, in which $D_{\text{match}}$ represents the count of documents where the decomposed structural backbone perfectly matches the ground truth. This constitutes a rigorous metric that requires zero structural errors.
@@
We compare our LingoEDU (built on Qwen3-4B (qwen3technicalreport)) against two categories of strong baselines. For all LLMs, we designed specific prompts to instruct them to output hierarchical JSON/Markdown structures: (1)Frontier LLMs: We evaluated SOTA models including GPT-4o/4.1 (gpt-4o), OpenAI o3/o4-mini (o3), Claude 3.7 Sonnet/4 (Claude4), DeepSeek-V3/R1 (deepseekai2025deepseekv3technicalreport), and Qwen3-235B (qwen3). All LLM results are averaged over three distinct runs. (2)Commercial Parsing APIs: We selected Jina Reader and Firecrawl, which are widely used for web-to-markdown conversion. We deployed test documents on a static server to allow URL-based access.
@@
Our method utilizes Qwen3-4B as the backbone. The training involved a two-stage process: (1) Continued Pre-training on $\sim$100k synthetic samples to learn layout patterns, followed by (2) Supervised Fine-Tuning (SFT) on thousands of meticulously manually annotated documents to align with human intent. All experiments were conducted on a Linux operating system running on a high-performance server equipped with an Intel Xeon 2.3GHz CPU, 1960GB of memory, and 8 NVIDIA A100 GPUs, each with 80 GB of VRAM.
3.1.2 Experiment Results
@@
Table 1 highlights the superiority of our explicit training paradigm. While top-tier commercial LLMs like Claude-4-Sonnet and OpenAI o3 achieve competitive structural scores (TED $\sim$5.1–5.5), other robust models such as DeepSeek-R1 and Qwen3 still struggle, plateauing at a TED of 6.2–9.9. Qualitative analysis reveals that general models often hallucinate non-existent sub-sections or flatten deep hierarchies to save generation tokens. Similarly, commercial parsing APIs lack semantic depth; Jina and Firecrawl exhibit high TED scores (>16>16) as they rely on shallow HTML tags and fail to capture implicit discourse structures found in complex PDFs. In contrast, our LingoEDU Decomposer demonstrates specialized efficiency by achieving a remarkable TED of 4.77 and a DLA of 49.60%, significantly outperforming the strongest baseline, Claude-4-Sonnet (+6.45% absolute DLA). This confirms that structural understanding requires dedicated supervision beyond what emergent prompting or reasoning models can provide.
@@
3.1.3 Ablation Studies
Table 2 validates the effectiveness of our design choices. First, the significant performance gap between “Indices Only” and our method (33.06% vs. 49.60% DLA) highlights that explicit text generation acts as a crucial semantic anchor for structural prediction. Second, the model exhibits remarkable data efficiency; even when scaled down to just 20% of the training data, it achieves a TED of 4.87 and retains over 91% of the full model’s accuracy.
Ablation Setting Variant TED DLA (%)
Output Formulation Indices Only 8.16 33.06
Indices + Text (Ours) 4.77 49.60
Data Scaling 20% Data 4.87 45.16
50% Data 4.85 48.79
100% Data 4.77 49.60
@@ 3.1.3 Ablation Studies
Finally, we investigate the impact of model scale on structural parsing capability, as shown in Table 3. Scaling the backbone from 1.7B to 4B yields clear improvements, reducing TED from 4.99 to 4.77. However, further scaling to 8B results in performance saturation: both TED and DLA slightly regress compared to the 4B model. This suggests that the 4B parameter range strikes an optimal balance for this specific task, whereas larger models may suffer from overfitting to the rigid output format without proportionally larger datasets.
   Model Size    TED ($\downarrow$)    DLA % ($\uparrow$)
   Qwen-1.7B    4.99    48.39
   Qwen-4B    4.77    49.60
   Qwen-8B    4.89    49.19
Table 3: Ablation study on backbone model scaling. Our 4B model achieves the best balance between structural error (TED) and relation accuracy (DLA).
Model / Method Multi-Doc QA Summarization Few-shot
HotpotQA 2Wiki Musique DuReader GovRep QMSum MultiN VCSum TREC Trivia SAMSum LSHT
Qwen3-8B 65.83 73.92 - - 26.85 19.60 - - 70.50 87.98 - -
Glyph 66.42 72.98 - - 25.53 19.78 - - 82.62 88.54 - -
Gemini-2.5-Pro
   Standard 35.20 38.10 28.55 7.15 4.10 15.80 4.05 5.80 46.50 59.85 20.45 26.10
   Self-Sum 37.78 39.90 30.77 7.79 4.34 16.53 4.44 6.17 49.00 62.31 21.89 29.50
   Ours (LingoEDU) 40.46 40.91 31.22 8.12 4.25 16.17 4.85 6.36 57.50 63.25 23.80 35.48
   $\Delta$ (vs. Standard) +14.94% +7.38% +9.35% +7.69% +2.44% +2.34% +19.75% +9.66% +23.66% +1.25% +11.39% +3.45%
GPT-4.1
   Standard 65.83 72.98 51.90 21.80 29.97 22.84 20.85 12.50 77.00 90.07 39.20 48.60
   Self-Sum 67.89 74.39 53.48 23.51 30.98 22.53 22.06 13.71 79.00 93.69 40.79 50.50
   Ours (LingoEDU) 70.11 74.68 54.86 25.34 31.56 23.30 23.50 14.62 80.00 93.76 41.68 52.50
   $\Delta$ (vs. Standard) +6.50% +2.33% +5.70% +16.24% +2.94% +0.61% +5.80% +8.96% +3.90% +4.10% +6.33% +8.02%
Table 4: Transposed results on LongBench. Datasets are grouped by task type (columns). $\Delta$ denotes the relative improvement of Ours over the Standard baseline for each specific backbone. Bold indicates best performance per backbone group; underlined indicates second-best.
@@ 3.2 Main Results on Downstream Long-Context Tasks
In this section, we address the second research question: Does structure-aware compression tangibly enhance performance for downstream tasks? We evaluate the utility of our EDU-based Context Compressor across two distinct scenarios: standard long-context benchmarks and complex open-domain Deep Search pipelines.
@@ 3.2.1 General Long-Context Understanding
We evaluate on LongBench (bai-etal-2024-longbench), a benchmark covering multi-document QA, summarization, and few-shot learning.
@@
We compare three configurations: (1) Standard, which feeds the full (or truncated) original context directly to the LLM; (2) Self-Sum, which utilizes the LLM itself to generate abstractive summaries of the context prior to processing; and (3) Glyph (cheng2025glyph), which applies an implicit compression baseline to serve as a reference for latent-based methods.
@@
Table 4 compares performance across Gemini-2.5-Pro and GPT-4.1 backbones. Overall, our EDU-based approach yields consistent improvements over the Standard baseline, with relative gains ($\Delta$) peaking at +23.66% on few-shot tasks. Regarding Multi-Document QA (HotpotQA, Musique), our method outperforms Self-Sum. We attribute this to the tendency of abstractive methods to lose critical entities required for multi-hop reasoning; in contrast, our explicit tree structure preserves precise evidence chains via original text indices, enabling the model to look up exact details without hallucination. As for Summarization Tasks, while Self-Sum is naturally strong, our method remains competitive (e.g., surpassing Self-Sum on MultiNews by +1.44 points). Notably, it significantly outperforms the external Glyph baseline, suggesting that retaining hierarchical structure is more effective than latent compression for information retention.
Ablation Studies on Node Ranking Strategies.
To isolate the impact of the ranking mechanism within LingoEDU, we decouple node selection from the reasoning backbone. We fix the generator as GPT-4.1 across all experiments and compare five configurations: (1) Standard: Feeds the full (or truncated) context directly without explicit filtering. (2) Random: Randomly selects nodes to match our compression budget, serving as a stochastic lower bound. (3) BM25: A sparse retrieval baseline relying on lexical overlap. (4) Self-Sum: Prompts the generator (GPT-4.1) itself to identify relevant nodes prior to reasoning. (5) Ours (Qwen3-Reranker): Semantically scores nodes using our lightweight Qwen3-Reranker-0.6B.
Table 5 demonstrates that our dedicated ranking approach consistently outperforms other strategies. While the Standard baseline is hindered by noise in long contexts, structured selection significantly boosts performance. Crucially, Ours surpasses Self-Sum across most datasets (e.g., +2.2% on HotpotQA, +1.83% on DuReader). This result underscores that a specialized, lightweight dense ranker (0.6B)—despite its size—offers superior evidence localization compared to the intrinsic selection capabilities of a general-purpose LLM, which often lacks the granularity for precise context pruning. Furthermore, the substantial margin over BM25 validates the necessity of semantic-aware filtering over surface-level matching.
Selection Strategy Multi-Doc QA Summarization Few-shot
HotpotQA 2Wiki Musique DuReader GovRep QMSum MultiN VCSum TREC Trivia SAMSum LSHT
Standard (No Selection) 65.83 72.98 51.90 21.80 29.97 22.84 20.85 12.50 77.00 90.07 39.20 48.60
Random 56.42 45.11 37.71 21.70 30.52 20.85 22.38 12.53 31.50 91.81 38.13 27.50
BM25 65.99 59.91 48.71 26.84 31.43 23.16 23.05 12.59 53.00 91.58 37.65 36.75
Self-Sum (LLM-Select) 67.89 74.39 53.48 23.51 30.98 22.53 22.06 13.71 79.00 93.69 40.79 50.50
Ours (Qwen3-Reranker 0.6B) 70.11 74.68 54.86 25.34 31.56 23.30 23.50 14.62 80.00 93.76 41.68 52.50
Table 5: Ablation Study on Ranking Models. All methods use GPT-4.1 as the final generator. We compare No Selection, Random, BM25, LLM Self-Selection, and Our Dedicated Reranker. Bold denotes the best performance.
Model Backbone HLE BrowseComp-ZH
DeepSeek-R1 9.0 9.5 13.6 +51.11% 18.7 19.4 20.4 +9.09%
Qwen3-235B-Thinking 14.2 14.7 15.5 +9.15% 8.7 9.0 12.8 +47.13%
DeepSeek-V3.1 14.5 14.8 15.6 +7.59% 29.1 29.8 38.8 +33.33%
DeepSeek-V3.2 20.0 20.6 21.2 +6.00% 31.1 32.2 34.6 +11.25%
Closed-Source Models
GPT-5 25.0 25.9 27.1 +8.40% 29.1 29.8 31.8 +9.28%
Claude Opus 4.1 14.0 14.8 15.5 +10.71% 20.8 21.5 23.2 +11.54%
Gemini 3 Pro (w/o Deep Think) 26.1 26.7 30.1 +15.33% 47.4 48.1 48.8 +2.95%
Table 6: Ablation study of the LingoEDU module on Deep Search benchmarks. Accuracy scores (%) are reported. Base: Standard Deep Search without compression; Self-Sum: Query-focused summarization; Ours (LingoEDU): Structural decomposition. $\Delta$ denotes the relative improvement of Ours over the Base baseline.
We further validate the effectiveness and efficiency of our framework through extensive supplementary analyses in Appendix C.
@@ 3.2.2 Impact on Deep Search
Unlike standard retrieval, Deep Search involves aggregating information from diverse, often noisy web sources to answer complex, open-ended queries. We integrated the LingoEDU module into a Deep Search pipeline and evaluated it on two challenging benchmarks(HLE (phan2025humanity) and BrowseComp-ZH (zhou2025browsecomp)) designed to push the limits of current LLMs.
@@
Deep search engines frequently ingest raw web pages cluttered with ads, navigation bars, and irrelevant links. As shown in Table 6, on the noise-intensive BrowseComp-ZH, our method delivers dramatic gains. Specifically, Qwen3-235B and DeepSeek-V3.1 show relative improvements of 47.13% and 33.33%, respectively. This confirms that our Decomposer acts as a semantic filter: by identifying logical EDUs, it effectively prunes “structural dead branches” while preserving the core content, a capability that standard summarization lacks in such high-entropy Chinese web contexts.
@@
Critically, our method remains additive even for the most advanced models like GPT-5 and Gemini 3 Pro (w/o Deep Think)(Table 6, rows 4-6). This indicates that even as model capacity grows, handling unstructured noise and long-context reasoning remain bottlenecks. Our explicit compression provides a complementary signal that enhances the internal processing of SOTA LLMs.